OpenStudy (anonymous):

A circle has a radius of 5cm . A sector of the circle has a central angle of 7pi/6 radians. Find the area of the sector. Do not round any intermediate computations. Round your answer to the nearest tenth. =cm^2

OpenStudy (anonymous):

\[\frac{ 7\pi }{ 6 }\]

OpenStudy (amistre64):

well, since 7pi/6 is some ration of 2pi ... the percentage of the total area of the circle is: 7pi/6 7/6 ----- = --- = 7/12 of the total circle area 2pi 2

OpenStudy (amistre64):

so ...\[A=\frac 7{12}\pi~r^2\]

OpenStudy (amistre64):

i tend to get confused with sector and segment ... so that might have to be modified if the area is not a slice of pizza shape

OpenStudy (amistre64):

 nah, its fine ...